Simplify or solve if it is an equation.
(-64)^1/3 =


(81)^1/4 =


SR(5x) * SR(11y)

Answer by stanbon(75887)

Simplify or solve if it is an equation.
(-64)^1/3 = -4
------------------

(81)^1/4 = 3
--------------------------

SR(5x) * SR(11y) = SR(55xy)
===========================
Cheers,
Stan H.